# SEFL: Enhancing Educational Assignment Feedback with LLM Agents

### v2

@@ 4.2 Multi-faceted Evaluation
We also evaluate the fine-tuned models’ output using a LLM-as-a-judge framework, a method gaining traction as a method for evaluating free text output (Liu et al. 2023; Zheng et al. 2024; Chen et al. 2023; Verga et al. 2024; Törnberg 2023; Naismith, Mulcaire, and Burstein 2023; Gilardi, Alizadeh, and Kubli 2023; Kocmi and Federmann 2023; Huang et al. 2024; Gu et al. 2024; Falk et al. 2025). The same 150 random instances are rated by the four LLMs, namely GPT-4o (Hurst et al. 2024), Claude3.5-Sonnet, Command-R+, and DeepSeek-V3 (Liu et al. 2024a). For the closed-source models’ LLM-as-a-judge experiments, we use their respective APIs and the total costs were approximately 10 USD.
@@
For every example the judge model receives the assignment prompt together with the two candidate feedback drafts. It is asked to decide which draft is better on the same four base criteria as the humans. The instruction forbids numeric grades or explanations and requires the judge to output exactly one character, A or B, producing a clean pairwise preference label. We report the full prompt in the supplementary material in the code repository.
5 Results
@@
In Figure 2, we show the win rates of models fine-tuned with SEFL vs. their non-tuned version, evaluated by both humans and LLM-based judges. A value above 50% indicates that the SEFL-tuned models are preferred over their original versions. We show an example of the feedback in Figure 4, where we depict the abridged prompt and feedback by two tuned an non-tuned models.
@@
Overall, human rater evaluations in Figure 2 (H#) show that the SEFL-tuned models often attain high win rates, surpassing 90% in several cases with respect to the smaller models. The human annotators differed in their views on the 8B model’s output quality; however, they generally converged on the observation that the fine-tuned 14B model produces superior feedback compared to its original version. By contrast, models not fine-tuned with SEFL had lower win rates, suggesting that SEFL provides an edge in generating more coherent and context-relevant feedback. In addition, we asked annotators whether the synthetic assignmentanswerfeedback sequences were consistent. In over 75% of cases, they confirmed the alignment between assignment, student response, and the feedback given, showing positive contextual relevance.
LLM-as-a-Judge Results.
For the LLM-as-a-judge evaluations (J#) in the same figure, we observe notable differences in win rates depending on the model and scale. The results largely mirror the human assessment trend up to the 3B scale. The results from the four LLM judges (J1: GPT-4o, J2: Claude-3.5-Sonnet, J3: Command-R+, J4: Deepseek-v3) reveal that SEFL-tuned models show varying levels of performance relative to their vanilla counterparts.333Models are picked based on their recency and performance on RewardBench (Lambert et al. 2024), JudgeBench (Tan et al. 2024), and JudgeArena (AtlaAI 2025). For instance, Qwen2.5-0.5B achieved the highest win rates across all four judges (62% on J3), indicating a consistent preference for the fine-tuned version. In contrast, larger models such as Llama-3.1-8B and Qwen2.5-14B exhibit lower win rates, particularly on J3 (16% and 10%, respectively), suggesting that fine-tuning with SEFL may yield diminishing returns or challenges at larger scales. The full judge prompt can be found in the supplementary material.
Agreement.
In Figure 3, we present the pairwise Cohen’s values (Cohen 1960) computed between each language-model judge and the human raters, to observe whether humans and LLM judges agree on which model gives better feedback. The agreement among humans was moderate to substantial: H1 and H3 reached , H1 and H2 , and H2 and H3 (Landis and Koch 1977). Among the models, Claude aligns most closely with both the other judges and the humans; DeepSeek follows, and GPT-4o shows the weakest match. Across all model and human pairs the coefficients range from to , which underlines the subjectivity of feedback evaluation. Command R+ sits at the lower extreme, returning values between and when compared with human raters and with other judges, which signals virtually no agreement. This signals that the human experts mainly agree on the quality of feedback, but for LLM judges this can differ. Overall, there are improvements to be made in terms of agreement between LLMs and humans in terms of feedback quality.
[…]
@@
| 14B-SEFL | “Clear and student friendly while retaining depth.” | “Uses short sentences and connects advice to rubric.” | “Groups comments by theme which avoids repetition.” |
@@
6 Discussion
6.1 Human Qualitative Insights
@@
In addition to the win rates in Figure 2, our human annotators provided rich qualitative feedback on the model outputs, which we show in Table 3. Generally, on the critical side, they noted that if a student answer is too short or incomplete, neither model explicitly flags the missing details. More specifically, Qwen2.5-0.5B was praised for clarity and concision, whereas Llama-3.2-3B tended to repeat assignment details without offering actionable guidance. Annotators observed that Llama-3.2-1B often gave more specific and constructive feedback but occasionally sounded harsh, while Llama-3.1-8B sometimes overlooked key aspects. Overall, although Qwen2.5-14B achieved high win rates (94, 77, 81 across three annotators), these insights suggest that even top-performing models could improve in error detection, tone refinement, and contextual sensitivity.
@@
To further quantify, in Figure 5, we plot for each model the net balance of optional rater comments in five qualitative categories: Actionability (AC), Goal Orientation (GO), User Friendliness (UF), Consistency (CO) and Student Autonomy (AY) (Carless et al. 2011; Wiggins 2012). Squares denote the base models and circles the SEFL-tuned variants. Horizontal whiskers give bootstrap Wilson 95% confidence intervals for the net balance. We compute these intervals on the proportion of positive remarks and then transform them to the net scale via , where is the total number of comments and the positive proportion. Annotators added comments only when a response stood out, so the plot reveals both the direction and the strength of impressions. We show several examples of annotations in Table 4. The annotators also had a choice if the feedback was non-nonsensical, which happened around 12% of the time mainly with the smaller sized models.
@@
The two smallest models, Qwen2.5-0.5B and Llama-3.2-1B, receive more negative than positive remarks on Consistency and Goal Orientation, matching earlier findings that they sometimes drift from the student answer or overlook core requirements. Llama-3.1-8B shows the only clearly positive balance in Actionability and User Friendliness, yet its interval for Consistency still lies below zero. Qwen2.5-14B gains more favorable notes on tone and clarity than the smaller models but still shows a negative alignment gap. With respect to Student Autonomy, there are mostly neutral comments.
| Method/Dataset | J1 | J2 | J3 | J4 |
| Book2Dial | 43 | 37 | 60 | 27 |
| SEFL | 57 | 63 | 40 | 73 |
6.2 LLM-as-a-Judge
@@
We used LLM judges to rate the feedback generated by SEFL-tuned models against their vanilla counterparts. This provides a rapid, scalable way to measure feedback quality, reducing the need for extensive human annotation. We let the LLM judges rate the same examples as the humans annotated. As shown in Figure 2, three out of four LLM judges consistently favored SEFL-tuned Qwen2.5-0.5B, Llama-3.2-1B, and Llama-3.2-3B. With Command-R, we notice that it performs worse than GPT-4o and Claude-3.5-Sonnet on JudgeArena. This indicates that the misalignment might have to do with instruction following capabilities of these models, this is highlighted by Command-R being weaker in Chatbot Arena (Zheng et al. 2024). Nonetheless, we see it as a practical first step for large-scale feedback comparisons in educational contexts. We recommend supplementing LLM-based assessments with targeted human evaluations for more granular insights, possibly aligning more with instructional objectives or possibly even fine-tuning LLMs with rubrics for better judges of long form text (Kim et al. 2024).
6.3 Comparison to Prior Work
@@
The work that is closest to ours is Book2Dial (Wang et al. 2024a). It is a framework that turns textbooks into synthetic conversations between a student model and a teacher model. The student only sees high-level cues such as section titles or key concepts, while the teacher has full access to the source passage, prompting a question-answer exchange that stays aligned with the book content. Instead, we focus on any type of assignment not limited to textbooks.
@@
To compare the two methods, we further fine-tune Qwen2.5-14B with the Book2Dial data in the same way as our SEFL-tuned version. We use the existing Book2Dial data consisting of 889 dialogues, we preprocess the data such that there is only a one-turn conversation, resulting in 5.3K conversation pairs. We then run the fine-tuned model over the same samples evaluated in Subsection 4.2 and evaluate them with the same 4 judges and evaluation criteria.444At the time of writing, we ran out of funding for the human annotators and thus compare here only with LLM-as-a-Judge.
Table 5 shows that three out of four judges prefer SEFL, yielding an average win rate of 58%. Only command-r-plus leans the other way with a 60% win rate for Book2Dial, but an average of 42% win rate. The finding is consistent with its earlier drift of seemingly failing to instruction follow. The results confirm that SEFL produces higher quality feedback than the textbook based dialogues in Book2Dial.
7 Conclusion
@@
We introduced SEFL, a framework that simulates teacherstudent interactions via two-agent LLMs to generate synthetic data for fine-tuning smaller models. Empirical and qualitative results show the potential of this framework for enhancing feedback in educational settings. This approach yields concise, context-sensitive feedback that often surpasses original instruction-tuned models under both LLM-as-a-judge and human evaluations. Yet human insights remain indispensable for capturing nuances like clarity and tone. SEFL provides a promising avenue for immediate, personalized feedback at scale, extending beyond the educational domain.
Limitations.
@@
We acknowledge that SEFL relies on synthetically generated assignments and errors, and are not real student submissions, which could have implications. Although this approach helps create large datasets, it risks producing feedback unaligned with authentic classroom contexts. Our evaluation also uses LLM-based judges, introducing potential biases related to each judge’s training data and objectives. Lastly, while we focused on short-answer tasks, longer or more domain-specific assignments may require specialized or more diverse synthetic data.

### v3

@@ 4.2.2. LLM-as-a-Judge
We also evaluate the fine-tuned models’ output using a LLM-as-a-judge framework, a method gaining traction for evaluating free text output Liu et al. (2023); Zheng et al. (2024); Chen et al. (2023); Verga et al. (2024); Törnberg (2023); Naismith et al. (2023); Gilardi et al. (2023); Kocmi and Federmann (2023); Huang et al. (2024); Gu et al. (2024); Falk et al. (2025). The three LLMs rate the same 450 random instances, namely GPT-4o Hurst et al. (2024), Claude3.5-Sonnet, and DeepSeek-V3 Liu et al. (2024a). For the LLM-as-a-judge experiments using closed-source models, we used the respective APIs, and the total cost was approximately 10 USD.
@@
For every example, the judge model receives the assignment prompt together with the two candidate feedback drafts. It is asked to decide which draft is better on the same four base criteria as the human annotators. The instruction forbids numeric grades or explanations and requires the judge to output exactly one character, A or B, producing a clean pairwise preference label.
5. Results
@@
In Figure˜2, we show the win rates of models fine-tuned with SEFL vs. their non-tuned version, evaluated by both humans and LLM-based judges. A value above 50% indicates that the SEFL-tuned models are preferred over their original versions. We show an example of the feedback in Figure˜4, where we depict the abridged prompt and feedback by a tuned and non-tuned model.
@@
Overall, human rater evaluations in Figure˜2 (H#) show that the SEFL-tuned models often achieve high win rates, surpassing 90% in several cases compared to the smaller models. The human annotators differed in their views on the 8B model’s output quality; however, they generally converged on the observation that the fine-tuned 14B model produces superior feedback compared to its original version. By contrast, models not fine-tuned with SEFL had lower win rates, suggesting that SEFL provides an edge in generating more coherent and context-relevant feedback. In addition, we asked annotators whether the synthetic assignmentanswerfeedback sequences were consistent. In over 75% of cases, they confirmed the alignment between assignment, student response, and the feedback given, showing positive contextual relevance.
LLM-as-a-Judge Evaluation.
For the LLM-as-a-judge evaluations (J#) in the same figure, we observe several differences in win rates depending on the model and scale. The results largely mirror the human assessment trend up to the 3B scale. The results from the three LLM judges (J1: GPT-4o, J2: Claude-3.5-Sonnet, J3: Deepseek-v3) reveal that SEFL-tuned models show varying levels of performance relative to their vanilla counterparts.444Models are picked based on their performance on RewardBench Lambert et al. (2024), JudgeBench Tan et al. (2024), and JudgeArena AtlaAI (2025). For instance, Qwen2.5-0.5B achieved the highest win rates across all three judges, indicating a consistent preference for the fine-tuned version. As the parameter size increases, the difference in win rate decreases (e.g., 58% on average for 8B). The full judge prompt will be released in the supplementary material.
Human and Model Agreements.
In Figure˜3, we present the pairwise Cohen’s values Cohen (1960) computed between each LLM judge and the human rater, to observe whether humans and LLM judges agree on which model gives better feedback. The agreement among humans was moderate to substantial: H1 and H3 reached , H1 and H2 , and H2 and H3 Landis and Koch (1977). Among the models, Claude aligns most closely with both the other judges and the humans; Deepseek follows, and GPT-4o shows the weakest match. Across all model and human pairs, the numbers range from 0.17 to 0.58, which shows the subjectivity of feedback evaluation. This suggests that human experts generally agree on the quality of feedback, but for LLM judges, this may vary. Overall, we see an opportunity for improvements in terms of agreement between LLMs and humans regarding feedback quality.
| Model | H1 sample | H2 sample | H3 sample |
@@
| 14B-SEFL | “Clear and student-friendly while retaining depth.” | “Uses short sentences and connects advice to the rubric.” | “Groups comments by theme, which avoids repetition.” |
@@
6. Discussion
6.1. Human Qualitative Insights
@@
In addition to the win rates in Figure˜2, our human annotators provided rich qualitative feedback on the model outputs, which we show in Table˜4. Generally, on the critical side, they noted that if a student’s answer is too short or incomplete, neither model explicitly flags the missing details. More specifically, Qwen2.5-0.5B was praised for clarity and concision, whereas Llama-3.2-3B tended to repeat assignment details without offering actionable guidance. Annotators observed that Llama-3.2-1B often gave more specific and constructive feedback but occasionally sounded harsh, while Llama-3.1-8B sometimes overlooked key aspects. Overall, although Qwen2.5-14B achieved high win rates (94, 77, 81 across three annotators), these insights suggest that even top-performing models could improve in error detection, tone refinement, and contextual sensitivity.
@@
To further quantify, in Figure˜5, we plot for each model the net balance of optional rater comments in five qualitative categories: Actionability (AC), Goal Orientation (GO), User Friendliness (UF), Consistency (CO), and Student Autonomy (AY) Carless et al. (2011); Wiggins (2012). Squares denote the base models and circles the SEFL-tuned variants. Horizontal whiskers provide 95% confidence intervals for the net balance using the bootstrap method. We compute these intervals on the proportion of positive remarks and then transform them to the net scale via , where is the total number of comments and the positive proportion. Annotators added comments only when a response stood out, so the plot reveals both the direction and the strength of impressions. We show several examples of annotations in Table˜5. The annotators also had the option to reject nonsensical feedback, which occurred around 12% of the time, primarily with the smaller-sized models.
@@
The two smallest models, Qwen2.5-0.5B and Llama-3.2-1B, receive more negative than positive remarks on Consistency and Goal Orientation, matching earlier findings that they sometimes drift from the student answer or overlook core requirements. Llama-3.1-8B shows the only clearly positive balance in Actionability and User Friendliness, yet its interval for Consistency still lies below zero. Qwen2.5-14B gains more favorable notes on tone and clarity than the smaller models, but still shows a negative alignment gap. Regarding Student Autonomy, there are mostly neutral comments. We show an example of the full output feedback of Qwen2.5-0.5B in Figure˜8 (Appendix˜C).
| Method/Dataset | J1 | J2 | J3 |
| Book2Dial | 43 | 37 | 27 |
| SEFL | 57 | 63 | 73 |
6.2. LLM-as-a-Judge
@@
We used LLM judges to rate the feedback generated by SEFL-tuned models against their untuned counterparts. This provides a scalable method for measuring feedback quality, thereby reducing the need for extensive human annotation. We let the LLM judges rate the same examples as the humans annotated. As shown in Figure˜2, 3 out of 3 LLM judges consistently favored SEFL-tuned Qwen2.5-0.5B, Llama-3.2-1B, and Llama-3.2-3B. We see it as a practical first step for large-scale feedback comparisons in educational contexts. We recommend supplementing LLM-based assessments with targeted human evaluations for more granular insights, possibly aligning more with instructional objectives or even fine-tuning LLMs with rubrics for better judgment of long-form text Kim et al. (2024).
6.3. Comparison to Prior Work
@@
The work that is most closely related to ours is Book2Dial Wang et al. (2024a). It is a framework that turns textbooks into synthetic conversations between a student model and a teacher model. The student only sees high-level cues, such as section titles or key concepts, while the teacher has full access to the source passage, prompting a question-and-answer exchange that remains aligned with the book content. Instead, we focus on any assignment, not limited to textbooks.
@@
To compare the two methods, we further fine-tune Qwen2.5-14B with the Book2Dial data in the same way as our SEFL-tuned version. We utilize the existing Book2Dial data, which consists of 889 dialogues, and preprocess the data to ensure that each conversation is one-turn, resulting in 5,300 conversation pairs. We then run the fine-tuned model over the same samples evaluated in Subsection 4.2 and assess them with the same 3 judges and evaluation criteria.555At the time of writing, we ran out of funding for the human annotators and thus compare here only with LLM-as-a-Judge. Table˜6 shows that 3 out of 3 judges prefer SEFL, yielding an average win rate of 64%. The results confirm that SEFL produces higher quality feedback than the textbook-based dialogues in Book2Dial.
7. Conclusion
@@
We introduced SEFL, a framework that simulates teacherstudent interactions via two-agent LLMs to generate synthetic data for fine-tuning smaller models. In this work, we aimed to explore how synthetic teacher-student interactions generated by LLMs can be leveraged for scalable feedback. We address this by demonstrating that these interactions, when generated through SEFL, can be used as high-quality fine-tuning data to enhance the feedback capabilities of smaller, more efficient language models. This approach yields concise, context-sensitive feedback that often surpasses original instruction-tuned models under both LLM-as-a-judge and human evaluations. Yet, human insights remain indispensable for capturing nuances such as clarity and tone. SEFL provides a promising avenue for immediate, personalized feedback at scale, extending beyond the educational domain.
Limitations
@@
We acknowledge that SEFL relies on short-form, synthetically generated assignments and errors, which are not real student submissions, and this could have implications. Although this approach helps create large datasets, it risks producing feedback unaligned with authentic classroom contexts. Our evaluation also utilizes LLM-based judges, which introduces potential biases related to each judge’s training data and objectives. Lastly, while we focused on short-answer tasks, longer or more domain-specific assignments may require specialized or more diverse synthetic data, which we leave to future work.